Divide Rs 300 between hena and rina in such a way that rina gets Rs 100 less than 3 times of what hena gets?

Answer by TimothyLamb(4379)

x = hena amount
y = rina amount
---
x + y = 300
y = 3x - 100
---
put the system of linear equations into standard form
---
x + y = 300
3x - y = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = hena amount = 100
y = rina amount = 200
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php